# Global Bifurcation of Non-Radial Solutions for Symmetric Sub-linear Elliptic Systems on the Planar Unit Disc

### v1

@@ 2. $(b)$
If there is an orbit type $(H)\in\mathfrak{M}_{1}$ with $\bar{\mathfrak{s}}:=\bar{\mathfrak{s}}(H)$ and for which $2\nmid|\mathfrak{J}(H)|$, then the system (1) admits an unbounded branch $\mathscr{C}$ of non-radial solutions with symmetries at least $({}^{\bar{\mathfrak{s}}}H)$. Moreover, we have the following alternative: there exists $M>0$ such that, either
@@
    for all $\alpha>-M$ with $(\alpha,0)\notin\Lambda^{\bm{K}}$ one has $\mathscr{C}\cap\{\alpha\}\times\mathscr{H}\not=\emptyset$.
@@
Notice that, for any critical point $(\alpha_{i},0)\in\mathfrak{J}(H)$, the numbers $\mathfrak{m}^{\bar{\mathfrak{s}}}(\alpha_{i}^{\pm},H)$ have the same parity (cf. Remark 4.5). Consider any other critical point $(\alpha_{j},0)\in\mathfrak{J}(H)$ with $\alpha_{i}<\alpha_{j}$. We will show that $\mathfrak{m}^{\bar{\mathfrak{s}}}(\alpha_{i}^{+},H)$ and $\mathfrak{m}^{\bar{\mathfrak{s}}}(\alpha_{j}^{-},H)$ are also of the same parity. Indeed, suppose for contradiction that $\mathfrak{m}^{\bar{\mathfrak{s}}}(\alpha_{i}^{+},H)$ and $\mathfrak{m}^{\bar{\mathfrak{s}}}(\alpha_{j}^{-},H)$ have different parities. Then there is a folding $m>\bar{\mathfrak{s}}$ such that the sets $\Sigma^{m}(\alpha_{i}^{+},H)$, $\Sigma^{m}(\alpha_{j}^{-},H)$ disagree for some odd numbers of indices (equivalently, the numbers $\mathfrak{n}^{m}(\alpha_{i}^{+},H)$, $\mathfrak{n}^{m}(\alpha_{j}^{-},H)$ have different parity). Consequently, there must be an intermediate critical point $(\alpha_{k},0)\in\mathfrak{J}(H)$ with $\alpha_{i}<\alpha_{k}<\alpha_{j}$ and $\mathfrak{i}^{m}(\alpha_{k},H)\neq 0$. This is in contradiction with the assumption of maximality for $\bar{\mathfrak{s}}$. It follows that the quantity $(-1)^{\mathfrak{m}^{\bar{\mathfrak{s}}}(\alpha_{*}^{-},H)}$ is constant for all critical points $(\alpha_{*},0)\in\mathfrak{J}(H)$. On the other hand, let $(\alpha_{k},0),(\alpha_{k+1},0)\in\mathfrak{J}(H)$ be any two consecutive critical points. From the definition of (38), the numbers $\mathfrak{i}^{\bar{\mathfrak{s}}}(\alpha_{k},H),\mathfrak{i}^{\bar{\mathfrak{s}}}(\alpha_{k+1},H)$ must be non-zero. With an argument similar to that the one used above, it can be shown that the numbers $\mathfrak{n}^{\bar{\mathfrak{s}}}(\alpha_{k}^{+},H)$ and $\mathfrak{n}^{\bar{\mathfrak{s}}}(\alpha_{k+1}^{-},H)$ have the same parity, from which it follows that
@@ 5. Motivating Example: Vibrating Membranes with Symmetric Coupling
satisfying the following assumptions:
@@
then the faces of the cube (see Figure LABEL:fig:6membranes) are naturally permuted as follows
(43) $\forall_{\sigma\in S_{4}\leq S_{6}}\;\;\sigma(u_{1},u_{2},u_{3},u_{4},u_{5},u_{6})^{T}=(u_{\sigma(1)},u_{\sigma(2)},u_{\sigma(3)},u_{\sigma(4)},u_{\sigma(5)},u_{\sigma(6)})^{T},\quad u\in V.$
@@
Under the assumptions ($E_{1}$)-($E_{3}$), the function $f$ clearly satisfies conditions ($A_{1}$)-($A_{3}$). For simplicity, let us also assume that $f$ satisfies the conditions ($A_{4}$) and ($A_{5}$) with the matrix $A(\alpha):=D_{u}f(0):V\rightarrow V$ given by
@@
The sigmoid function introduces a nonlinear dependence of the coupling in (41) on the bifurcation parameter $\alpha\in\mathbb{R}$. Moreover, the asymptotic behaviour of (46) imposes coupling strength saturation limits which might reflect physical constraints such as, for example, bounded ranges for certain material densities.
@@
If $c,d\in\mathbb{R}$ are such that $c+4d,c-2d,c\neq 0$, then (42) satisfies condition ($B_{3}$). Moreover, together with the strict monotonicity of the eigenvalues (47) and the distinctness of the numbers $s_{nm}$, it follows that each critical point of (41) can be specified by a unique triple $(n,m,j)\in\Sigma$ with the notation
@@
Let us make the system (41) more specific by choosing the following values for the constants $a,c,d\in\mathbb{R}$
@@
Notice that the critical set for (41) consists of exactly five critical points that can be expressed, using the notation (48), as follows
@@
In order to employ the results presented in previous sections to the bifurcation problem (41), we must first identify the maximal orbit types in $\Phi_{0}(G;\mathscr{H}_{1}\setminus\{0\})$. The following G.A.P code can be used to generate a complete list of the sets $\mathfrak{M}_{1,j}$ and the corresponding Basic Degrees $\deg_{\mathcal{V}_{1,j}}$ for $j=1,2,3$.
@@ 28 View(max_orbtyps);
The output of the above program can be expressed using amalgamated notation and an arbitrary $m>0$ folding (cf. (23) and (24)) as follows:
@@
Take $j\in\{0,2,3\}$ and let $(H)$ be any orbit type in $\mathfrak{M}_{1,j}$. From the coefficients of the Basic Degrees $\deg_{\mathcal{V}_{1,j}}\in A(G)$, it is clear that the fixed point space $\mathcal{V}_{1,j}^{H}$ has odd dimension. Since the critical set (50) is manageably small, we can compute the cardinalities (26) for each of the critical points $(\alpha_{n,m,j},0)\in\Lambda$ to obtain the quantities
@@
Next, we can determine the global behaviour of some of the branches predicted using Theorem (4.1) by establishing the parity of the set (38). Notice that $\max\limits_{(\alpha,0)\in\Lambda}\{\mathfrak{s}(\alpha,H)\}\nmid 2$ is obtained for exactly three critical points $(\alpha_{1,3,0},0),(\alpha_{1,3,2},0),(\alpha_{1,3,3},0)$, i.e.
@@
Therefore, any branch of non-trivial solutions with symmetries at least $({}^{3}H)$ bifurcating from $(\alpha_{1,3,0},0),(\alpha_{1,3,2},0),(\alpha_{1,3,3},0)$ consists only of non-radial solutions and is unbounded. Of course, since the critical set (50) is finite, one needs only sum the local bifurcation invariants (LABEL:) to arrive at the same conclusion directly from the Rabinowitz alternative (cf. Theorems 4.2 and 4.3).

### v2

@@ 2. $(b)$
If there is an orbit type $(H)\in\mathfrak{M}_{1}$ for which $2\nmid|\mathfrak{J}(H)|$, then the system (1) admits an unbounded branch $\mathscr{C}$ of non-radial solutions with symmetries at least $({}^{\bar{\mathfrak{s}}}H)$, where $\bar{\mathfrak{s}}:=\bar{\mathfrak{s}}(H)$. Moreover, one has the following alternative: there exists $M>0$ such that, either
@@
    for all $\alpha<-M$ with $(\alpha,0)\notin\Lambda^{\bm{K}}$ one has $\mathscr{C}\cap\{\alpha\}\times\mathscr{H}\not=\emptyset$.
@@
Notice that, for any critical point $(\alpha_{i},0)\in\mathfrak{J}(H)$, the numbers $\mathfrak{m}^{\bar{\mathfrak{s}}}(\alpha_{i}^{\pm},H)$ have the same parity (cf. Remark 4.5). Without loss of generality, assume that $|\mathfrak{J}(H)|>1$ and consider any other critical point $(\alpha_{j},0)\in\mathfrak{J}(H)$ with $\alpha_{i}<\alpha_{j}$. We will show that $\mathfrak{m}^{\bar{\mathfrak{s}}}(\alpha_{i}^{+},H)$ and $\mathfrak{m}^{\bar{\mathfrak{s}}}(\alpha_{j}^{-},H)$ are also of the same parity. Indeed, suppose for contradiction that $\mathfrak{m}^{\bar{\mathfrak{s}}}(\alpha_{i}^{+},H)$ and $\mathfrak{m}^{\bar{\mathfrak{s}}}(\alpha_{j}^{-},H)$ have different parities. Then there is a folding $m>\bar{\mathfrak{s}}$ such that the sets $\Sigma^{m}(\alpha_{i}^{+},H)$, $\Sigma^{m}(\alpha_{j}^{-},H)$ disagree for some odd numbers of indices (equivalently, the numbers $\mathfrak{n}^{m}(\alpha_{i}^{+},H)$, $\mathfrak{n}^{m}(\alpha_{j}^{-},H)$ have different parity). Consequently, there must be an intermediate critical point $(\alpha_{k},0)\in\mathfrak{J}(H)$ with $\alpha_{i}<\alpha_{k}<\alpha_{j}$ and $\mathfrak{i}^{m}(\alpha_{k},H)\neq 0$. However, this is in contradiction with the assumption of maximality for $\bar{\mathfrak{s}}$. It follows that the quantity $(-1)^{\mathfrak{m}^{\bar{\mathfrak{s}}}(\alpha_{*}^{-},H)}$ is constant for all critical points $(\alpha_{*},0)\in\mathfrak{J}(H)$. On the other hand, let $(\alpha_{k},0),(\alpha_{k+1},0)\in\mathfrak{J}(H)$ be any two consecutive critical points. From the definition of (39), the numbers $\mathfrak{i}^{\bar{\mathfrak{s}}}(\alpha_{k},H),\mathfrak{i}^{\bar{\mathfrak{s}}}(\alpha_{k+1},H)$ must be non-zero. With an argument similar to that the one used above, it can be shown that the numbers $\mathfrak{n}^{\bar{\mathfrak{s}}}(\alpha_{k}^{+},H)$ and $\mathfrak{n}^{\bar{\mathfrak{s}}}(\alpha_{k+1}^{-},H)$ have the same parity, such that
@@ 5. Motivating Example: Vibrating Membranes with Symmetric Coupling
For compatibility with the results derived in previous sections, we assume that (43) satisfies the following conditions:
@@
then the faces of the cube (see Figure 1) are naturally permuted as follows
(44) $\sigma(u_{1},u_{2},u_{3},u_{4},u_{5},u_{6})^{T}:=(u_{\sigma(1)},u_{\sigma(2)},u_{\sigma(3)},u_{\sigma(4)},u_{\sigma(5)},u_{\sigma(6)})^{T},\quad\forall_{\sigma\in S_{4}\leq S_{6}},\;u\in V.$
@@
Under the assumptions ($E_{1}$)-($E_{3}$), the function $f$ clearly satisfies conditions ($A_{1}$)-($A_{3}$). Let us also assume that $f$ satisfies the conditions ($A_{4}$) and ($A_{5}$) with the matrix $A(\alpha):=D_{u}f(0):V\rightarrow V$ given by
@@
The sigmoid function introduces a nonlinear dependence of the coupling in (42) on the bifurcation parameter $\alpha\in\mathbb{R}$. Moreover, the asymptotic behaviour of (47) can be understood as the imposition of coupling strength saturation limits, which might reflect physical constraints such as, for example, bounded ranges for certain material densities.
@@
Notice also that, if $c,d\in\mathbb{R}$ are such that $c+4d,c-2d,c\neq 0$, then the critical set associated with (43) is discrete. Together with the strict monotonicity of the eigenvalues (48) and the distinctness of the numbers $s_{nm}$, it follows that each critical point of (42) can be specified by a unique triple $(n,m,j)\in\Sigma$ with the notation
@@
Let us specify a particular bifurcation problem associated with system (42) by choosing the following values for the constants $a,c,d\in\mathbb{R}$
@@
Notice that the critical set for (42) consists of exactly five critical points that can be distinguished, using the notation (49), as follows
@@
In order to employ the results presented in previous sections to our bifurcation problem, we must first identify the maximal orbit types in $\Phi_{0}(G;\mathscr{H}_{1}\setminus\{0\})$. The following G.A.P code can be used to generate a complete list of the sets $\mathfrak{M}_{1,j}$ and the corresponding Basic Degrees $\deg_{\mathcal{V}_{1,j}}$ for $j=1,2,3$.
@@ 28 View(max_orbtyps);
For any $m>0$ and $j\in\{0,2,3\}$, the output of the above program can be used to describe the the set of maximal orbit types $\mathfrak{M}_{m,j}\subset\Phi_{0}(G)$ the basic degrees $\deg_{\mathcal{V}_{m,j}}\in A(G)$ (cf. (24) and (25)), as follows:
@@
We can deduce, from the coefficients of the Basic Degrees $\deg_{\mathcal{V}_{1,j}}\in A(G)$ (cf. Remark 4.3), that the fixed point space $\mathcal{V}_{1,j}^{H}$ has odd dimension for any $j\in\{0,2,3\}$ and $(H)\in\mathfrak{M}_{1,j}$. Since the critical set (51) is manageably small, we can manually compute the cardinalities (27) for each of the critical points $(\alpha_{n,m,j},0)\in\Lambda$ to obtain the quantities
@@
Next, we can determine the global behaviour of some of the branches predicted using Theorem (4.1) by establishing the parity of the set (39). Notice that $2\nmid\max\limits_{(\alpha,0)\in\Lambda}\{\mathfrak{s}(\alpha,H)\}$ is obtained for exactly three critical points $(\alpha_{1,3,0},0),(\alpha_{1,3,2},0),(\alpha_{1,3,3},0)$, i.e.
@@
Therefore, any branch of non-trivial solutions with symmetries at least $({}^{3}H)$ bifurcating from $(\alpha_{1,3,0},0),(\alpha_{1,3,2},0),(\alpha_{1,3,3},0)$ consists only of non-radial solutions and is unbounded. Of course, since the critical set (51) is finite, one needs only sum the relevant local bifurcation invariants to arrive at the same conclusion directly from the Rabinowitz alternative (cf. Theorems 4.2 and 4.3).